SPB/SFX 
	- are working on achieving a good focus @ 6 keV, which currently is not as good as expected.
 
	- Digitizers are ready, have been tested by CAS ==> to be be installed / integrated probably next week
 
	- Different GUIs were reporting varying values for STATUS on different clients / machines ==> ARound will make sure a proper description of the issue (which machines, which clients, when, ...) will be brought to the attention of the proper people @ CAS.
 
 
FXE 
	- The Karabo FAI went well, FXE thanks explicitly for the support received.
 
 
SQS 
	- The pnCCD is being set up. AE will be required to help for axes commisisoning next week. ==> this should happen today.
 
 
SCS 
	- MTeichmann attended, fresh from vacation, nothing to report, everything is "awesome".
 
 
 MID 
	- Commissioning the cryomanipulator. Had to use a couple of work arounds in order to align it with the beam (travel range issues it appears).
 
	- The AGIPD Julabo chiller failed on Saturday, it has to be sent in. Meetings this mornings should clarify what other options remain to support the user experiments as planned.
 
 
HED 
	- User experiment went very well.
 
	- Currently commissioning ion pumps in the optics hutch
 
	- Starting on Wednesday loop 6 is to be commisioned (exp. hutch components) together with AE
 
 
OPT Lasers 
XRO 
XPD 
VAC 
DET 
	- Question to ITDM: Did the crash of git on Saturday affect the calibration database (there were apparently issues @ FXE for 2 hours) ==> No. If there would have been problems, they would not have resolved with nobody doing anything, the system requires manual intervention. So most probably the problem was somewhere else...
 
 
AE 
	- SCS had a problem with the fast ADC, was fixed Friday. Did this gave further problems during the weekend ==> MTeichmann to check.
 
 
ITDM 
EET 
CAS 
SRP 
TS 
AoB 
 |